# SVFAP: Self-supervised Video Facial Affect Perceiver

### v1

@@ 3.1.4 Decoder
where $\mathbf{X}_{i}$ is the output of stage $i$ ($i\in\{2,3\}$) in the encoder, $k^{i-1}$ is the upscaling factor. Moreover, to facilitate the decoder to be aware of spatiotemporal features in different levels, summation-based multi-scale fusion is employed to aggregate the outputs of three stages in the encoder, i.e.,
@@ 3.2 Downstream Fine-tuning
Based upon the pre-trained encoder, we apply global average pooling to the extracted spatiotemporal feature and append a fully-connected network for final prediction. We use different loss functions for different types of video-based facial affect analysis tasks. For the classification task, the cross-entropy loss is employed, i.e.,
@@
where $\hat{\mathbf{y}}$ denotes the prediction, $\mathbf{y}$ is the target, and $K$ is the number of classes. While for the regression task, we compute the mean square error between the prediction and the target:
$\mathcal{L}_{\textrm{reg}}=||\mathbf{y}-\hat{\mathbf{y}}||^{2}$ (13)
@@ 4.1 Implementation Details
TPSBT architecture. To meet different needs in real-world applications, we build two versions (i.e., base and small) of TPSBT as the encoder of SVFAP. For the small version TPSBT-S, $C=384$, $M_{1}=8$, $M_{2}=4$, $M_{3}=2$. For the base version TPSBT-B, $C=512$, $M_{1}=12$, $M_{2}=6$, $M_{3}=3$. The model size and computational cost of the small version is approximately half of the base version. Besides, for both versions, the spatial bottleneck token number $G$ is set to 8.
Self-supervised pre-training. We pre-train TPSBT on a large-scale audio-visual dataset of human speech, VoxCeleb2 [62]. It includes more than 1 million video clips for over 6K celebrities extracted from about 150K videos uploaded to YouTube. We use its development set for self-supervised pre-training, which has 1,092,009 video clips from 145,569 videos. The original video is with a resolution of $224\times 224$. Since the speaker face typically does not cover the whole screen, we only use $160\times 160$ patch in the upper center of each video frame to filter irrelevant background information and reduce the computational cost. We sample 16 frames from each video clip using a temporal stride of 4, resulting in $8\times 10\times 10$ input tokens after patch embedding when using the patch size of $2\times 16\times 16$.
We conduct experiments using the PyTorch framework with 4 Nvidia GeForce RTX 3090 GPUs. For the hyperparameter setting, we mainly follow VideoMAE [23]. Specifically, we adopt an AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The base learning rate is $43e-$ 4 and the weight decay is 0.05. The overall batch size is 256. We linearly scale the base learning rate with respect to the overall batch size, i.e., lr=base learning rate×batch $\textrm{lr}=\textrm{base learning rate}\times\frac{\textrm{batch size}}{256}$. Besides, we employ a cosine decay learning rate scheduler. By default, we pre-train the model for 100 epochs with 5 warmup epochs and it takes about 5-6 days in our setting.
Downstream fine-tuning. The input video clip size is also $16\times 160\times 160$ and the temporal stride is 4 in most cases. We adopt an AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The base learning rate is $31e-$ 3 and the overall batch size is 96. Other hyperparameters are basically same as those in pre-training and can also refer to [23] for more details. We fine-tune the pre-trained model for 100 epochs with 5 warmup epochs. For inference, we uniformly sample two clips along the temporal axis for each video sample and then compute the average score as the final prediction.
@@ 4.2.1 Datasets
DFEW consists of 16,372 video clips which are extracted from more than 1,500 high-definition movies. Each video clip is annotated by 10 well-trained annotators with 7 basic emotions (i.e., anger, disgust, fear, happy, sad, surprise, and neutral). To align with previous studies [5, 12], we evaluate the proposed method on 11,697 single labeled clips using the default 5-fold cross-validation protocol.
@@
eNTERFACE05 is also an audio-visual emotion dataset which contains about 1,200 video clips from more than 40 subjects. Each subject is asked to simulate six emotions, including anger, disgust, fear, happy, sad, and surprise. To make a fair comparison with previous work [8], we employ a 5-fold subject-independent cross-validation protocol.
@@ 4.2.2 Ablation Studies
Training schedule. We then explore the effect of training schedule length. As shown in Fig. 4, we find that as the pre-training process goes on, the pre-training loss decreases steadily and the downstream fine-tuning performance improves consistently. This finding is in line with previous works on self-supervised learning [22, 24, 23]. It also should be noted that we do not observe clear performance saturation when reaching default maximum epochs (i.e., 100), which indicates that longer pre-training could further improve model performance in downstream fine-tuning. However, due to limited computational resources, we leave it for future work and hope the community can conduct follow-up study.
Moreover, we visualize several reconstructed video samples using our best pre-trained model in Fig. 5. Note that these samples are randomly selected from the test set of VoxCeleb2 thus can not be seen by our model during pre-training. We find that even under such a high masking ratio (i.e., 90%), SVFAP still can generate satisfying reconstruction results, especially for dynamic facial expressions. This indicates that, benefited from the challenging masked facial video autoencoding task, our model can reason over high-level and meaningful spatiotemporal semantics from limited visible input to recover masked information.
Dataset scale. We investigate how the model behaves when pre-trained with different dataset scales. For this purpose, we randomly generate a range of subsets with increasing sizes from the whole dataset, i.e., 1%, 5%, 10%, and 20%. Note that we proportionally increase the pre-training epochs for these subsets to ensure the same pre-training cost. The results are shown in Table I. We can observe that larger pre-training data scale leads to consistently better fine-tuning results. This is expected as more diversified training samples typically result in better generalization. It is also worth noting that, even with 11K unlabeled data, our method still achieves promising performance, which demonstrates that our SVFAP is a data-efficient self-supervised video facial affect learner.
@@
TABLE II: Ablation study of model variants. TP: Temporal Pyramid. SBT: Spatial Bottleneck Transformer. FLOPs-P: Pre-training FLOPs.
@@
Model variants. We ablate different model variants of TPSBT, including 1) no TP and SBT, i.e., the vanilla ViT baseline, by removing temporal pyramid learning modules and replacing SBT with a standard Transformer of similar size in the last two stages. 2) only TP, by replacing SBT with a standard Transformer of similar size. 3) only SBT, by removing temporal pyramid learning modules. As presented in Table II, we have the following observations: 1) When compared with the vanilla ViT baseline, TP largely reduces computational overhead (about 30% and 14% FLOPs reduction for fine-tuning and pre-training respectively) and achieves comparable performance, with only 1.1M additional parameters. 2) SBT not only reduces large computations (about 35% FLOPs and 11% FLOPs-P reduction) but also achieves the best performance among all variants. 3) When combines TP and SBT together, our default full model achieves the lowest computational cost while almost maintaining the best performance of SBT. Specifically, TPSBT significantly reduces about 43% and 15% FLOPs during fine-tuning and pre-training and outperforms the vanilla ViT baseline by 0.92% UAR and 0.40% WAR, with the sacrifice of a slight increase of model parameters (1.2M). These results indicate that temporal pyramid learning and spatial bottleneck mechanism can greatly remove redundant spatiotemporal information in 3D facial videos (from temporal and spatial perspectives respectively) and help the model to concentrate on informative one, thus contributing to lower computational cost and better model performance.
Multi-scale fusion. We explore the effect of multi-scale spatiotemporal feature fusion during both self-supervised pre-training and downstream fine-tuning. As shown in Table III, we find that employing multi-scale fusion during pre-training achieves better results than those that do not use it, which shows that integrating spatiotemporal features in different levels for decoder reconstruction can help the encoder to learn more useful representations. However, we observe slight performance degradation when using it during fine-tuning. This result could be partly ascribed to the simple and parameter-free temporal upsampling method (i.e., nearest-neighbor interpolation) used for multi-scale fusion. Therefore, we only use the high-level feature in the last stage of the encoder during downstream fine-tuning.
TABLE III: Ablation study of multi-scale fusion.
@@
TABLE IV: Ablation study of spatial bottleneck tokens. FLOPs-P: Pre-training FLOPs.
@@
Spatial bottleneck tokens. SBT utilizes spatial attention to generate several global semantic bottleneck tokens for spatial redundancy elimination and computational cost reduction, thus it is necessary to investigate how the number of spatial bottleneck tokens influence model performance. Table IV presents the ablation results. We see that too few tokens (i.e., 4) hurt the model performance as it might be too aggressive to perform spatial compression and incur critical information loss. Besides, too many tokens (i.e., 16) increase model computational cost but do not contribute to significantly better results. Therefore, we set the spatial bottleneck token number to 8 by default.
TABLE V: Ablation study of temporal downsampling. FLOPs-P: Pre-training FLOPs.
@@
Temporal downsampling. We explore the effect of three different methods for temporal downsampling in temporal pyramid learning, including simple average pooling, max pooling, and the default strided convolution. As presented in Tab. V, we find that the convolution-based method outperforms the other two pooling methods with the sacrifice of slightly more model parameters and negligible FLOPs increase, which justifies our default design choice.
TABLE VI: Comparison with state-of-the-art supervised and self-supervised pre-trained models on split 1 of DFEW. SSL: Self-Supervised Learning or not.
@@
Video Swin [59] $\times$ ImageNet-1K+K400 Image+Video Swin-S 50 55 57.13 69.96
Video Swin [59] $\times$ ImageNet-21K+K400 Image+Video Swin-B 88 93 59.38 71.90
@@
TABLE VII: Comparison with state-of-the-art methods on DFEW.
@@
TABLE VIII: Comparison with state-of-the-art methods on FERV39k.
@@
TABLE IX: Comparison with state-of-the-art methods on MAFW. AN: Anger. DI: Disgust. FE: Fear. HA: Happiness. NE: Neutral. SA: Sadness. SU: Surprise. CO: Contempt. AX: Anxiety. HL: Helplessness. DS: Disappointment.
@@ 4.2.3 Comparison with Previous Pre-trained Models
In this section, we show the effectiveness of the proposed self-supervised learning method by comparing it with previous state-of-the-art supervised and self-supervised pre-trained models. The supervised models contain four advanced video Transformers pre-trained on large-scale labeled video or image datasets, including TimeSformer [56], MViT [57], MViTv2 [58], and Video Swin Transformer [59]. The self-supervised part involves eight cutting-edge models pre-trained on massive unlabeled data (most are facial images or videos), including FAb-Net [47], TCAE [48], FaceCycle [49], BMVC’20 [50], MoCo [69], SwAV [51], VideoMAE [23], and FaRL [52]. It should be noted, all self-supervised models except VideoMAE are 2D models, i.e., they can not process video inputs directly. Therefore, we add a standard Transformer block on top of them to enable temporal sequential modeling. Besides, all models share the same evaluation protocol to ensure a fair comparison.
The comprehensive comparison results on the split 1 of DFEW are presented in Table VI. From the table, we have the following key observations:
@@
  • Compared with self-supervised pre-trained models, SVFAP also achieves the best performance. Notably, when using the same Voxceleb2 dataset for pre-training, both our base and small model improves FAb-Net, TCAE, FaceCycle, and BMVC’20, by a significant margin, which indicates that masked facial video autoencoding is an effective task for self-supervised pre-training. Our method also beats two contrastive learning-based methods (i.e., MoCo and SwAV), verifying the advantage of generative paradigm in self-supervised learning. When compared to VideoMAE, SVFAP-S still shows better performance while having 2.9$\times$ fewer parameters and 4.5$\times$ fewer FLOPs. Finally, our method also outperforms the best-performing FaRL (i.e., 3.72% UAR and 2.66% WAR improvements for SVFAP-B, 0.32% UAR and 0.80% WAR improvements for SVFAP-S), although FaRL has much more FLOPs and parameters and is pre-trained on a huge multimodal dataset. To summarize, the above comparison results with self-supervised models show that our SVFAP is an effective and efficient self-supervised video facial affect perceiver.
In addition to the evaluation in full data regime, we also conduct experiments to investigate the generalization ability of SVFAP under few-shot settings. To this end, we randomly select 50%, 20%, and 10% samples from the training set of DFEW split 1 to obtain a series of new training sets while keeping the test set unchanged. For simplicity, we only choose several representative models in Table VI for evaluation. Note that, for the method with more than one model, we only use the best one. The results are reported in Fig. 6. We can observe that: 1) As expected, the performance of all methods degrades accordingly when fewer training samples are used. 2) SVFAP consistently outperforms all compared methods under each few-shot setting, which verifies the strong adaptation ability of the proposed methods in the low data regime. This ability is particularly important considering the longstanding data scarcity issue in affective computing. Notably, even with 10% (about 935 samples) training data, our method still achieves promising results (more than 50% UAR and 63% WAR).

### v2

@@ 3.1.4 Decoder
where $\mathbf{X}_{i}$ is the output of stage $i$ ($i\in\{2,3\}$) in the encoder, $k^{i-1}$ is the upscaling factor. Moreover, to enable the decoder to be aware of spatiotemporal features at different levels, summation-based multi-scale fusion is employed to aggregate the outputs of three stages in the encoder, i.e.,
@@ 3.2 Downstream Fine-tuning
Based upon the pre-trained encoder, we apply global average pooling to the extracted spatiotemporal feature and append a fully connected network for final prediction. We use different loss functions for different types of video-based facial affect analysis tasks. For the classification task, the cross-entropy loss is employed, i.e.,
@@
where $\hat{\mathbf{y}}\in\mathbb{R}^{K}$ denotes the prediction, $\mathbf{y}\in\mathbb{R}^{K}$ is the target, and $K$ is the number of emotion classes. For the regression task, we compute the mean square error between the prediction and the target:
$\mathcal{L}_{\textrm{reg}}=\frac{1}{K}||\mathbf{y}-\hat{\mathbf{y}}||^{2}$ (13)
where $K$ is the number of emotion dimensions.
@@ 4.1 Implementation Details
TPSBT architecture. To meet different needs in real-world applications, we build two versions (i.e., base and small) of TPSBT as the encoder of SVFAP. For the small version TPSBT-S, $C=384$, $M_{1}=8$, $M_{2}=4$, $M_{3}=2$. For the base version TPSBT-B, $C=512$, $M_{1}=12$, $M_{2}=6$, $M_{3}=3$. The model size and computational costs of the small version are approximately half of the base version. Besides, for both versions, the spatial bottleneck token number $G$ is set to 8.
Self-supervised pre-training. We pre-train TPSBT on a large-scale audio-visual dataset of human speech, VoxCeleb2 [63]. It includes more than 1 million video clips for over 6K celebrities extracted from about 150K videos uploaded to YouTube. We use its development set for self-supervised pre-training, which has 1,092,009 video clips from 145,569 videos. As shown in Fig. 3, the resolution of the original videos in VoxCeleb2 is $224\times 224$, where faces typically appear in the upper-central part of the video. The remaining parts display the shoulders and neck, along with extraneous background information. Therefore, we only use a 160x160 patch from the upper-central location of the video, allowing us to remove irrelevant information while also reducing the model’s input size, thus lowering computational costs. We sample 16 frames from each video clip using a temporal stride of 4, resulting in $8\times 10\times 10$ input tokens after patch embedding when using the patch size of $2\times 16\times 16$.
Figure 3: The illustration of the face patch location for VoxCeleb2 videos.
We conduct experiments using the PyTorch framework with 4 Nvidia GeForce RTX 3090 GPUs. For the hyperparameter setting, we mainly follow VideoMAE [24]. The main differences include the learning rate, batch size, and training epochs. Specifically, we adopt an AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The base learning rate is $43e-$ 4 and the weight decay is 0.05. The overall batch size is 256. We linearly scale the base learning rate with respect to the overall batch size, i.e., lr=base learning rate×batch $\textrm{lr}=\textrm{base learning rate}\times\frac{\textrm{batch size}}{256}$. Besides, we employ a cosine decay learning rate scheduler. By default, we pre-train the model for 100 epochs with 5 warmup epochs and it takes about 5-6 days in our setting.
Downstream fine-tuning. The input video clip size is also $16\times 160\times 160$ and the temporal stride is 4 in most cases. We adopt an AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The base learning rate is $31e-$ 3 and the overall batch size is 96. Other hyperparameters are the basically same as those in pre-training and can also refer to [24] for more details. We fine-tune the pre-trained model for 100 epochs with 5 warmup epochs. For inference, we uniformly sample two clips along the temporal axis for each video sample and then compute the average score as the final prediction.
@@ 4.2.1 Datasets
DFEW consists of 16,372 video clips which are extracted from more than 1,500 high-definition movies. Each video clip is annotated by 10 well-trained annotators with 7 basic emotions (i.e., anger, disgust, fear, happy, sad, surprise, and neutral). To align with previous studies [6, 13], we evaluate the proposed method on 11,697 single-labeled clips using the default 5-fold cross-validation protocol.
@@
eNTERFACE05 is also an audio-visual emotion dataset that contains about 1,200 video clips from more than 40 subjects. Each subject is asked to simulate six emotions, including anger, disgust, fear, happy, sad, and surprise. To make a fair comparison with previous work [9], we employ a 5-fold subject-independent cross-validation protocol.
@@ 4.2.2 Ablation Studies
Training schedule. We then explore the effect of training schedule length. As shown in Fig. 5, we find that as the pre-training process goes on, the pre-training loss decreases steadily and the downstream fine-tuning performance improves consistently. This finding is in line with previous works on self-supervised learning [23, 25, 24]. It also should be noted that we do not observe clear performance saturation when reaching default maximum epochs (i.e., 100), which indicates that longer pre-training could further improve model performance in downstream fine-tuning. However, due to limited computational resources, we leave it for future work and hope the community can conduct follow-up studies.
Moreover, we visualize several reconstructed video samples using our best pre-trained model in Fig. 6. Note that these samples are randomly selected from the test set of VoxCeleb2, whose speakers are disjoint with those from the development set used for pre-training. Thus, they are not seen by our model during pre-training. We find that even under such a high masking ratio (i.e., 90%), SVFAP still can generate satisfying reconstruction results, especially for dynamic facial expressions. This indicates that, benefiting from the challenging masked facial video autoencoding task, our model can reason over high-level and meaningful spatiotemporal semantics from limited visible input to recover masked information.
Dataset scale. We investigate how the model behaves when pre-trained with different dataset scales. For this purpose, we randomly generate a range of subsets with increasing sizes from the whole dataset, i.e., 1%, 5%, 10%, and 20%. Note that we proportionally increase the pre-training epochs for these subsets to ensure the same pre-training cost. The results are shown in Table I. We can observe that larger pre-training datasets generally lead to better fine-tuning results. This is expected as more diversified training samples typically result in better generalization. It is also worth noting that, even with 11K unlabeled data, our method still achieves promising performance, which demonstrates that our SVFAP is a data-efficient self-supervised video facial affect learner.
@@
TABLE II: Ablation study of the masking ratio. FLOPs-P: Pre-training FLOPs.
Percentage
#Params
FLOPs
FLOPs-P
UAR WAR
75% 77.6 43.6 18.4 60.89 73.91
85% 77.6 43.6 14.7 61.87 74.77
90% 77.6 43.6 12.9 62.63 74.81
95% 77.6 43.6 11.2 59.93 73.17
Masking ratio. The influence of the masking ratio is presented in Table II. We can find that the masking ratio of 90% has the best performance. The lower masking ratios of 75% and 85% achieve worse performance, although the encoder accepts more video tokens as input and has higher computational costs during pre-training. A higher masking ratio of 95% results in lower pre-training cost, however, the performance degrades significantly. These results are consistent with previous findings [24, 25]. Therefore, we set 90% as the default masking ratio.
TABLE III: Ablation study of model variants. TP: Temporal Pyramid. SBT: Spatial Bottleneck Transformer. FLOPs-P: Pre-training FLOPs.
@@
Model variants. We ablate different model variants of TPSBT, including 1) no TP and SBT, i.e., the vanilla ViT baseline, by removing temporal pyramid learning modules and replacing SBT with a standard Transformer of similar size in the last two stages. 2) only TP, by replacing SBT with a standard Transformer of similar size. 3) only SBT, by removing temporal pyramid learning modules. As presented in Table III, we have the following observations: 1) When compared with the vanilla ViT baseline, TP largely reduces computational overhead (about 30% and 14% FLOPs reduction for fine-tuning and pre-training respectively) and achieves comparable performance, with only 1.1M additional parameters. 2) SBT not only reduces large computations (about 35% FLOPs and 11% FLOPs-P reduction) but also achieves the best performance among all variants. 3) When combining TP and SBT together, our default full model achieves the lowest computational costs while almost maintaining the best performance of SBT. Specifically, TPSBT significantly reduces about 43% and 15% FLOPs during fine-tuning and pre-training and outperforms the vanilla ViT baseline by 0.92% UAR and 0.40% WAR, with the sacrifice of a slight increase of model parameters (1.2M). These results indicate that temporal pyramid learning and spatial bottleneck mechanism can greatly remove redundant spatiotemporal information in 3D facial videos (from temporal and spatial perspectives respectively) and help the model to concentrate on the informative one, thus contributing to lower computational costs and better model performance.
Multi-scale fusion. We explore the effect of multi-scale spatiotemporal feature fusion during both self-supervised pre-training and downstream fine-tuning. As shown in Table IV, we find that employing multi-scale fusion during pre-training achieves better results than those that do not use it, which shows that integrating spatiotemporal features in different levels for decoder reconstruction can help the encoder to learn more useful representations. However, we observe slight performance degradation when using it during fine-tuning. This result could be partly ascribed to the simple and parameter-free temporal upsampling method (i.e., nearest-neighbor interpolation) used for multi-scale fusion. Another reason is that, compared to low-level details, high-level emotional semantics are more crucial for downstream affect analysis tasks. Therefore, we only use the high-level feature from the last stage of the encoder during fine-tuning.
TABLE IV: Ablation study of multi-scale fusion.
@@
TABLE V: Ablation study of spatial bottleneck tokens. FLOPs-P: Pre-training FLOPs.
@@
Spatial bottleneck tokens. SBT utilizes spatial attention to generate several global semantic bottleneck tokens for spatial redundancy elimination and computational cost reduction, thus it is necessary to investigate how the number of spatial bottleneck tokens influence model performance. Table V presents the ablation results. We see that too few tokens (i.e., 4) hurt the model performance as it might be too aggressive to perform spatial compression and incur critical information loss. Besides, too many tokens (i.e., 16) increase model computational costs but do not contribute to significantly better results. Therefore, we set the spatial bottleneck token number to 8 by default.
TABLE VI: Ablation study of temporal downsampling. FLOPs-P: Pre-training FLOPs.
@@
Temporal downsampling. We explore the effect of three different methods for temporal downsampling in temporal pyramid learning, including simple average pooling, max pooling, and the default strided convolution. As presented in Tab. VI, we find that the convolution-based method outperforms the other two pooling methods with the sacrifice of slightly more model parameters and negligible FLOPs increase, which justifies our default design choice.
TABLE VII: Comparison with state-of-the-art supervised and self-supervised pre-trained models on split 1 of DFEW. SSL: Self-Supervised Learning or not.
@@
Video Swin [60] $\times$ ImageNet-1K+Kinetics-400 Image+Video Swin-S 50 55 57.13 69.96
Video Swin [60] $\times$ ImageNet-21K+Kinetics-400 Image+Video Swin-B 88 93 59.38 71.90
@@
$\rho$BYOL [71] Kinetics-400 Video SlowOnly-R50 32 43 58.60 69.81
SVT [72] Kinetics-400 Video TimeSformer 121 198 57.07 70.01
@@
TABLE VIII: Comparison with state-of-the-art methods on DFEW.
@@
TABLE IX: Comparison with state-of-the-art methods on FERV39k.
@@
TABLE X: Comparison with state-of-the-art methods on MAFW. AN: Anger. DI: Disgust. FE: Fear. HA: Happiness. NE: Neutral. SA: Sadness. SU: Surprise. CO: Contempt. AX: Anxiety. HL: Helplessness. DS: Disappointment.
@@ 4.2.3 Comparison with Previous Pre-trained Models
In this section, we show the effectiveness of the proposed self-supervised learning method by comparing it with previous state-of-the-art supervised and self-supervised pre-trained models. The supervised models contain four advanced video Transformers pre-trained on large-scale labeled video or image datasets, including TimeSformer [57], MViT [58], MViTv2 [59], and Video Swin Transformer [60]. The self-supervised part involves eight cutting-edge models pre-trained on massive unlabeled data (most are facial images or videos), including FAb-Net [48], TCAE [49], FaceCycle [50], BMVC’20 [51], MoCo [70], SwAV [52], $\rho$BYOL [71], SVT [72], VideoMAE [24], and FaRL [53]. It should be noted, all self-supervised models except $\rho$BYOL, SVT, and VideoMAE are 2D models, i.e., they can not process video inputs directly. Therefore, we add a standard Transformer block on top of them to enable temporal sequential modeling. Besides, all models share the same evaluation protocol to ensure a fair comparison.
The comprehensive comparison results on the split 1 of DFEW are presented in Table VII. From the table, we have the following key observations:
@@
  • Compared with self-supervised pre-trained models, SVFAP also achieves the best performance. Notably, when using the same Voxceleb2 dataset for pre-training, both our base and small model improve FAb-Net, TCAE, FaceCycle, and BMVC’20, by a significant margin, which indicates that masked facial video autoencoding is an effective task for self-supervised pre-training. Our method also beats contrastive learning-based methods (i.e., MoCo, SwAV, and SVT), verifying the advantage of the generative paradigm in self-supervised learning. When compared to VideoMAE, SVFAP-S still shows better performance while having 2.9$\times$ fewer parameters and 4.5$\times$ fewer FLOPs. Finally, our method also outperforms the best-performing FaRL (i.e., 3.72% UAR and 2.66% WAR improvements for SVFAP-B, 0.32% UAR and 0.80% WAR improvements for SVFAP-S), although FaRL has much more FLOPs and parameters and is pre-trained on a huge multimodal dataset. To summarize, the above comparison results with self-supervised models show that our SVFAP is an effective and efficient self-supervised video facial affect perceiver.
In addition to the evaluation in full data regime, we also conduct experiments to investigate the generalization ability of SVFAP under few-shot settings. To this end, we randomly select 50%, 20%, and 10% samples from the training set of DFEW split 1 to obtain a series of new training sets while keeping the test set unchanged. For simplicity, we only choose several representative models in Table VII for evaluation. Note that, for the method with more than one model, we only use the best one. The results are reported in Fig. 7. We can observe that: 1) As expected, the performance of all methods degrades accordingly when fewer training samples are used. 2) SVFAP consistently outperforms all compared methods under each few-shot setting, which verifies the strong adaptation ability of the proposed methods in the low data regime. This ability is particularly important considering the longstanding data scarcity issue in affective computing. Notably, even with 10% (about 935 samples) training data, our method still achieves promising results (more than 50% UAR and 63% WAR).